OpenStudy (anonymous):

Find the slope of the lines possing through the points, then decided whether the lines are parallel, perpendicular or intersect (1,-1), (-4,-4),(1,1),(4,-4)

OpenStudy (aroub):

(1,-1) (-4,-4) (1,1)(4,-4) ok so (1,-1)(-4,-4) / (1,1)(4,-4) their s -4-1/4-1=-5/3 lopes are perpendicular -4+1/-4-1 =3/5